If a, b and c are the roots of the equation x


1
Expert's answer
2021-08-30T06:18:42-0400

Question: If a, b, and c are the roots of the equation "x\u00b3-4x\u00b2+3x+5", what is the value of (a+1)(b+1)(c+1)?


Answer:

for cubic equations, "ax\u00b3+bx\u00b2+cx+d", with roots a,b and c

"a+b+c = -b/a\\

ab+bc+ac = c/a\\

abc = -d/a"


"(a+1)(b+1)(c+1) = (a+b+c) + (ab+bc+ac) + abc + 1 = 4 + 3 -5 + 1 = 3"